Let (X,(tua)) be a space and ((X) infinity, (tau) infinity) its one point compactification .prove that ((X)infinity, (tau)infinity) is compact
1
Expert's answer
2021-07-20T17:52:46-0400

Let "\{U_i \sub X_\infty\}_{i\in I}" be an open cover. We need to show that this has a finite subcover.

Since we have an open cover, "\exist i_\infty \in I" such that "U_{i_\infty}" is an open neighbourhood of "\infty". But the only open subsets containing "\infty" are of the form "(X \setminus C) \cup \{\infty\}", so "\exist" a compact closed subset "C" of "X" such that "X \setminus C \sub U_{i_\infty}".


Also, "\{U_i \sub X\}_{i\in I}" is an open cover of "C". Since "C" is compact, "\exist" a finite subcover "\{U_i \sub X\}_{i\in J}" where "J \sub I".


Hence we have that "\{U_i⊂X\}_{i∈J}∪U_{i_∞}" is a finite subcover of the original cover. Thus our one-point compactification is compact.